OpenStudy (anonymous):

According to genetic theory, every plant of a particular species has a 25% chance of being red-flowering, independently of all other plants. Among 10 plants of this species, what is the chance that fewer than 4 are red-flowering? My answer was 28431/1048576. I did it like this: The chance that there is 1 red and 9 others: 1/4 * (3/4)^9 The chance that there are 2 red and 8 others: 1/4 * 1/4 * (3/4)^8 The chance that there are 3 red and 7 others: 1/4 * 1/4 * 1/4 *(3/4)^7. All the others added up made 28431/1048576. But this is wrong. Anybody could give me a hint in the right directi

OpenStudy (anonymous):

it is wrong because you also need the number of ways to get each out come

OpenStudy (anonymous):

for example one red and 9 others has nine ways it can happen

OpenStudy (anonymous):

what you need is \[P(x=k)=\binom{10}{k}(\frac{1}{4})^k(\frac{3}{4})^{n-k}\] you are missing the \(\binom{10}{k}\) part

OpenStudy (anonymous):

actually i meant \[P(x=k)=\binom{10}{k}(\frac{1}{4})^k(\frac{3}{4})^{10-k}\]

OpenStudy (anonymous):

ahh the number of combinations ... stupid me! I will give it a shot

OpenStudy (anonymous):

The results would be: 754515/1048576

OpenStudy (anonymous):

The answer is not correct hmmm

OpenStudy (anonymous):

Take this as bonus my friend : the answer is 0.775875.

OpenStudy (anonymous):

you need the details ???? consult your binomial theory. you see the mean=np=10*(0.25) then it goes on and on..............lol

OpenStudy (anonymous):

binomial distribution: p=0.25 n=10 X<=3